OpenStudy (anonymous):

Use the method of cylindrical shells to find the volume V generated by rotating the region bounded by the given curves about the y-axis. y = x^2 y=0 x=1

OpenStudy (amistre64):

HA!! shells, i knew they were comin

OpenStudy (amistre64):

since we is going around the y axis; its simpler right now

OpenStudy (anonymous):

yup :( i never used this method before ><

OpenStudy (amistre64):

the radius of each shell moves from x=0 to x=1 right?

OpenStudy (anonymous):

yup

OpenStudy (amistre64):

and the height of each shell is determined by: f(x)

OpenStudy (anonymous):

wait, i know why its x=1, but why is it x=0

OpenStudy (amistre64):

so the area of each shell, when flattened out is height* width...

OpenStudy (amistre64):

umm... y axis means x=0 they are the same thing

OpenStudy (anonymous):

ohh okay gotcha

OpenStudy (amistre64):

the height = f(x) the width = 2pi [x] the area of any given shell = 2pi x[f(x)] so integrate that from [a,b] = [0,1] in this case

OpenStudy (amistre64):

2pi [S] x(x^2) dx ; [0,1] 2pi x^4 ------- = (pi x^4)/4 = pi/4 4

OpenStudy (anonymous):

how did you get rid of the 4th power in the numerator?

OpenStudy (amistre64):

you dont... why would you?

OpenStudy (anonymous):

but you put (pi x^4)/4 = pi/4 where did the 4th power go?

OpenStudy (amistre64):

F(0) = 0 F(1) = pi/4 F(1) - F(0) = pi/4

OpenStudy (amistre64):

over 2 lol

OpenStudy (amistre64):

i forgot i changed 2/4 to 1/2.... doh!

OpenStudy (amistre64):

\[\frac{2\pi x^4}{4} = \frac{\pi x^4}{2}\]

OpenStudy (anonymous):

much clearer, lol thank you :D

OpenStudy (amistre64):

when im wrong; say im wrong...not ask me how im right lol. I just assume im right ;)

OpenStudy (anonymous):

lol will do xD

OpenStudy (amistre64):

you understand the mechanics of the shell method tho?

OpenStudy (anonymous):

all the methods like the washer, disk, etc all those methods seem so similiar to eachother and i get them all confused. because in the end youre always taking the antiderivative and plugging in the boundaries right?

OpenStudy (anonymous):

like i understand theres certain formulas you have to follow for each method, which i have to memerize lol all of them seem really similiar

OpenStudy (anonymous):

so for the shell method, basically youre taking half of it, and spreading it out which is why it turns to be a rectangle shape?

OpenStudy (amistre64):

when we flatten out the "shell" we get a flat rectangular piece that the area is eay to determine right?

OpenStudy (anonymous):

correct

OpenStudy (amistre64):

the width = 2pi x...can you tell me why?

OpenStudy (amistre64):

we aint using half the shell, but the whole thing

OpenStudy (anonymous):

i think it has to do with the radius, is that why the width is 2 pi x?

OpenStudy (anonymous):

ohh is it the circumference of the cylnder?

OpenStudy (amistre64):

it does, but i want to make sure you understand this :) its very basic and easy...exactly lol

OpenStudy (anonymous):

lol i took a careful look at the cylinder and realized haha xD

OpenStudy (amistre64):

you did good :)

OpenStudy (amistre64):

the shell method makes things easier for some problems

OpenStudy (anonymous):

thanks ;) so basically for the shell method we're plugging in f(x) into the formula and taking the antiderivative of it? then plug in the boundaries and solve?

OpenStudy (amistre64):

lets think that thruough; we want to add up all the areas...which is what intagrating does; each area = 2pi x[f(x)]

OpenStudy (amistre64):

{S} 2pi x[f(x)] dx is what we do right?

OpenStudy (amistre64):

if there is a constant we can pull it aside right?

OpenStudy (anonymous):

mhmm

OpenStudy (anonymous):

the 2pi

OpenStudy (amistre64):

2 pi goes out and what are we left with inside?

OpenStudy (anonymous):

x*f(x)

OpenStudy (amistre64):

excatly; so we integrate 'x*f(x)'

OpenStudy (amistre64):

do you se how that differes from your first statement of ; we int f(x)?

OpenStudy (anonymous):

yesss, thanks for clearing that up for me :) this method seems a bit easier than the others lol

OpenStudy (amistre64):

look at the disc method

OpenStudy (anonymous):

oh, i have one question. when you solve for an integral, is it possible for it to be negative? or does it always have to be positive

OpenStudy (amistre64):

the disc method take all the areas of a given circle; with a radius of f(x) and adds them up

OpenStudy (amistre64):

volume is always a positive number

OpenStudy (amistre64):

what is the formula for the the area of a circle ?

OpenStudy (anonymous):

uh oh. darn i should of took the absolute value. i was so sure i did it right too. lol but thank you :)

OpenStudy (anonymous):

and area of circle is pi r^2

OpenStudy (amistre64):

so when we add up all the areas of the cicles thru integration we do: {S} pi [f(x)]^2 dx right? adding up the areas of circles

OpenStudy (anonymous):

so f(x) is the radius?

OpenStudy (amistre64):

look at the drawing i did in this one

OpenStudy (amistre64):

and tell me what the radius of each given circle is?

OpenStudy (anonymous):

oh, f(x) xD haha

OpenStudy (amistre64):

that is all these volume of rotation problems amopunt to :)

OpenStudy (anonymous):

i remember there was this one method i think where you had to divide by two, is there a method that involves that? or am i making stuff up lol :p

OpenStudy (amistre64):

i dont think thats a volume of rotation one :) in some area propblems we are asked to find the area under a curve that hopes the y axis... and if the symmetry is the same from left to right side of the y axis; we get a skewed result

OpenStudy (amistre64):

lets try something simple and see; suppose we want to find the area of a square that is sitting halfway between the sides

OpenStudy (amistre64):

lets say 4 high and 4 wide; the area should be 16 right?

OpenStudy (anonymous):

correct

OpenStudy (amistre64):

if we integrate f(x) = y = 4; from -2 to 2 we know we should get 16 right?

OpenStudy (anonymous):

how come its from -2 and 2?

OpenStudy (amistre64):

{S} 4 dx -> 4x 4(2) - 4(-2) = 8 -- 8 = 16, so that one doesnt apply to the /2 thing you discussed lol

OpenStudy (amistre64):

the distance from -2 to 2 = 4 right?

OpenStudy (anonymous):

yupp

OpenStudy (amistre64):

and a square with sides = 4 would straddle the y axis would sit from -2 to 2 on the x axis right?

OpenStudy (anonymous):

correct. i gotcha now :) lol

OpenStudy (anonymous):

if the prob asked to rotate around the x-axis, its generally the same idea of solving it like we did with the y axis right?

OpenStudy (amistre64):

i pic is worth a thousand words :)

OpenStudy (anonymous):

it really is. thank you for that :D

OpenStudy (amistre64):

yes, but we just need to make sure we are rotating it properly; make sure the numbers come out correctly

OpenStudy (amistre64):

once we can draw a picture, the rest is just intuitive

OpenStudy (anonymous):

gotcha. i shall try and attempt the rest of these shell method questions. thank you so so much for all your help. :D you make it possible for me to understand calculus lol xD thank you times infinity !

OpenStudy (amistre64):

youre welcome :)